OpenStudy (anonymous):

A 970 kg motorboat accelerates away from a dock at 2.7 m/s2. Its propeller provides a thrust force of 3.9 kN. What is the drag force exerted by the water on the boat?

OpenStudy (anonymous):

F=ma m=970kg a=2.7 m/s^2 F=(970)(2.7) F=2619N Since the motorboat is accelerating AWAY from the dock, the thrust force will be negative. *Don't forget to convert 3.9kN to N Drag force=F_1+F_2=-3900N+2619N Answer: -1281N or -1.218kN

OpenStudy (anonymous):

I too agree with the above solution...its explicitly based upon the newtons dynamic equation..net force=mass times acceleration.. or thrust force-drag force (here viscous force)=mass times acceleration.

OpenStudy (theraggedydoctor):

;PPP